Calculate the change in entropy of gases in the following cases: a) A 3.0 mol sample of an ideal gas expands reversibly and isothermally at 350 K until its volume doubled. b) The temperature of 1.0 mol of an ideal monatomic gas is raised reversibly from 200 K to 300 K, with its volume kept constant.
1
Expert's answer
2020-01-03T09:13:18-0500

The change of entropy of an ideal gas


"\Delta S(T,V)=n C_V \ln\frac{T_f}{T_i}+nR\ln\frac{V_f}{V_i}"

(a) "T_f=T_i", "V_f=2V_i"


"\Delta S(T,V)=nR\ln\frac{V_f}{V_i}=3.0\times 8.3\ln 2=17.3\:\rm J/K"

(b) "T_i=200\:{\rm K}, \; T_f=300\:{\rm K}," "V_f=V_i"


"\Delta S(T,V)=n C_V \ln\frac{T_f}{T_i}=1.0\times 12.5\times\ln\frac{300}{200}=5.0\:\rm J/K"